OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (texaschic101):

o = one step equations t = two step equations o + t = 1070 ---> t = 1070 - o 3o - 2t = 110 now sub 1070 - o in for t in the 2nd equation 3o - 2t = 110 3o - 2(1070 - o) = 110 3o - 2140 + 2o = 110 5o - 2140 = 110 5o = 110 + 2140 5o = 2250 o = 450 now sub 450 in for o in the 1st equation... o + t = 1070 450 + t = 1070 t = 1070 - 450 t = 620 450 one step equations auditioned today.